Biochem.

With the average amount of protein in the U.S. diet consisting of approximately 90 g, what amount of niacin can be synthesized?

Assume here that tryptophan makes up 1% of protein by weight.
Answer #1

Tryptophan makes up 1% of protein = 0.01 * 90 = 0.9g = 900mg.

It is observed that on the molar basis 60mg of tryptophan is equivalent to 1 mg niacin.

So 1 gm of tryptophan is equivalent to 1/60mg of niacin.

Therefore 900mg of tryptophan is equivalent to 1/60*900 = 15mg of niacin.

15mg of niacin can be synthesized.